# Does Editing Provide Evidence for Localization?thanks: Blogpost version to appear in ICLR Blogposts 2025

### v1

@@ Abstract
A basic aspiration for interpretability research in large language models is to ”localize” semantically meaningful behaviors to particular components within the LLM. There are various heuristics for finding candidate locations within the LLM. Once a candidate localization is found, it can be assessed by editing the internal representations at the corresponding localization and checking whether this induces model behavior that is consistent with the semantic interpretion of the localization. The question we address here is: how strong is the evidence provided by such edits? To assess localization, we want to assess the effect of the optimal intervention at a particular location. The key new technical tool is a way of adapting LLM alignment techniques to find such optimal localized edits. With this tool in hand, we give an example where the edit-based evidence for localization appears strong, but where localization clearly fails. Indeed, we find that optimal edits at random localizations can be as effective as aligning the full model. In aggregate, our results suggest that merely observing that localized edits induce targeted changes in behavior provides little to no evidence that these locations actually encode the target behavior.
@@ 1 Introduction
A basic goal of interpretability research for large language models is to map semantically meaningful behavior to particular subcomponents of the model. Semantically meaningful encompasses a wide range of things, e.g., “when asked for directions to the Eiffle tower, the model gives directions to Paris”, “the model responds truthfully”, or “the model will refuse to response”. The aim is to find, e.g., neurons, circuits, or regions of representation space that control these behaviors. If we could find such localizations, we could use them as building blocks to understand complex model behaviors. Many interpretability approaches can be understood in terms of the following idealized template [[, e.g.,]]zou2023representation, arditi2024refusal, wang2023backdoor, chen2024truth, wei2024assessing, li2024inference, meng2022locating, vig2020causal, geiger2021causal, soulos2019discovering, finlayson2021causal, wang2022interpretability, chan2022causal, hanna2024does, conmy2023towards, todd2023function, hendel2023context:
@@ 2 Background and results from ITI
$\bm{r}^{l+1}$ $:=\bm{r}^{l}+W^{l}\bm{o}=\bm{r}^{l}+\sum_{h=1}^{H}W_{h}\bm{o}_{h}\in\mathbb{R}^{DH}$ (4)
where $r^{l}\in\mathbb{R}^{DH}$ is the residual stream before layer $l$, $\text{Attn}_{h}^{l}$ is the $h$-th attention module at layer $l$, with $\bm{o}_{h}^{l}$ being its output. $\bm{o}^{l}$ is the concatenated head outputs. $W^{l}$ is the project-out matrix, that applies $H$ independent linear transformations to the corresponding head outputs. Finally $\bm{r}^{l+1}$ is residual stream output after layer $l$.
@@ 3 Editing Localized Heads Modifies the Output as Expected
We find that interventions at the localized heads are more effective than interventions at random heads. In fig. 1(a) we report the Info*Truth score (average truthfulness score times average informativeness score). We find that using localized heads have significantly higher Truth*Info scores than using random heads (p-value $1.6\times 10^{-8}$). In fact, using random heads often doesn’t have noticeable effect on the truthfulness at all, as shown in fig. 1(b), fig. 1(c) .
@@
Figure 1: Localized heads perform much better than random when using ITI interventions. We observe better Truth*Info scores, better truth-info score tradeoff, as well as better MC-KL tradeoff.
This appears to add further evidence that the localized heads are “special” for the truthfulness concept. However, this strong association could be because the intervention and localization are “correlated”, since both use statistics of the same activations (determined by the design of the data, etc). E.g. for heads with very low probing accuracy, the estimated intervention vectors could be very noisy, and thus the interventions could be less effective.
@@ 4 Finding “optimal” interventions
This formulation connects weight edits to representation edits. However, it doesn’t yet allow us to localize edits to specific heads — while $\bm{\theta}^{l}$ can be read as concatenation of headwise intervention vectors, the projected $W^{l}\bm{\theta}^{l}$ have no corresponding interpretions. Therefore, we can’t restrict the edits to specific heads by imposing structure on $\bm{b}^{l}$’s.
@@ 6 Intervening a single head is just as effective
It is now clear that edit-based evidence does not provide strong evidence for localization in the 16 head setup. However, a possible way of saving localization would be to argue that 16 heads is too many, giving too much leeway to induce any behavior we want with editing. For example, if we edited half the heads of the model, it would not be surprising if we could make the model do anything we wanted. Accordingly, we might hope that there is still a valid syllogism of the form “the localized edit is extremely constrainted” and “edits at this location optimally control the target behavior” implies “the target behavior is localized to this location”.
To test this, we now focus on the single head case. The procedure is simple: we randomly sample $24$ single heads, one at a time, and search for optimal interventions. The distribution of the best Truth*Info scores is shown in fig. 4(a). We find $5$ single-heads that are as effective fig. 4(b), and none of them has high probing accuracy. Notice that, still, none of these heads can be understood as localizing the truthfulness concept. The reason is that there are multiple distinct locations that work equally well! That is, even in the extreme case of a very localized edit that replicates the target behavior essentially optimally, we still cannot conclude that there is evidence supporting localization.
@@ 7 Are the Probing-Localized Heads Anything Special?
So far what we mean by localization, is that we can change model generation on target concept by an edits at this location. And our experiments show no evidence for this type of localization, and probing-localized heads play no special role.
@@
It’s important to note that the model’s probabilities for fixed responses, do not directly correspond to what the model actually generates. Even if the model assigns a higher probability to a truthful response than an untruthful one, it may still not generate the truthful response if the fixed dataset is off-policy (i.e. both probabilities are low). This highlights the well-known gap between what a model ”knows” (which is the motivation behind probing) and what it ultimately generates [joshi2023personas, wang2020language, kadavath2022language, saunders2022self, burns2022discovering].

### v2

# Does Editing Provide Evidence for Localization? [^1] Blogpost version to appear in ICLR Blogposts 2025

@@ Abstract
A basic aspiration for interpretability research in large language models is to “localize” semantically meaningful behaviors to particular components within the LLM. There are various heuristics for finding candidate locations within the LLM. Once a candidate localization is found, it can be assessed by editing the internal representations at the corresponding localization and checking whether this induces model behavior that is consistent with the semantic interpretation of the localization. The question we address here is: how strong is the evidence provided by such edits? To evaluate the localization claim, we want to assess the effect of the optimal intervention at a particular location. The key new technical tool is a way of adapting LLM alignment techniques to find such optimal localized edits. With this tool in hand, we give an example where the edit-based evidence for localization appears strong, but where localization clearly fails. Indeed, we find that optimal edits at random localizations can be as effective as aligning the full model. In aggregate, our results suggest that merely observing that localized edits induce targeted changes in behavior provides little to no evidence that these locations actually encode the target behavior.
@@ 1 Introduction
A basic goal of interpretability research for large language models is to map semantically meaningful behavior to particular subcomponents of the model. Semantically meaningful encompasses a wide range of things, e.g., “when asked for directions to the Eiffel Tower, the model gives directions to Paris”, “the model responds truthfully”, or “the model will refuse to respond”. The aim is to find, e.g., neurons, circuits, or regions of representation space that control these behaviors. If we could find such localizations, we could use them as building blocks to understand complex model behaviors. Many interpretability approaches can be understood in terms of the following idealized template [[, e.g.,]]zou2023representation, arditi2024refusal, wang2023backdoor, chen2024truth, wei2024assessing, li2024inference, meng2022locating, vig2020causal, geiger2021causal, soulos2019discovering, finlayson2021causal, wang2022interpretability, chan2022causal, hanna2024does, conmy2023towards, todd2023function, hendel2023context:
@@ 2 Background and results from ITI
$\bm{r}^{l+1}$ $:=\bm{r}^{l}+W^{l}\bm{o}=\bm{r}^{l}+\sum_{h=1}^{H}W_{h}^{l}\bm{o}_{h}\in\mathbb{R}^{DH}$ (4)
where $r^{l}\in\mathbb{R}^{DH}$ is the residual stream before layer $l$, $\text{Attn}_{h}^{l}$ is the $h$-th attention module at layer $l$, with $\bm{o}_{h}^{l}$ being its output. $\bm{o}^{l}$ is the concatenated head outputs. $W^{l}$ is the project-out matrix, that applies $H$ independent linear transformations to the corresponding head outputs. Finally, $\bm{r}^{l+1}$ is residual stream output after layer $l$.
@@ 3 Editing Localized Heads Modifies the Output as Expected
We find that interventions at the localized heads are more effective than interventions at random heads. In fig. 1(a) we report the Info*Truth score (average truthfulness score times average informativeness score). We find that using localized heads have significantly higher Info*Truth scores than using random heads (p-value $1.6\times 10^{-8}$). In fact, using random heads often doesn’t have noticeable effect on the truthfulness at all, as shown in fig. 1(b), fig. 1(c).
@@
Figure 1: Localized heads perform much better than random when using ITI interventions. We observe better Info*Truth scores, better truth-info score tradeoff, as well as better MC-KL tradeoff.
This appears to add further evidence that the localized heads are “special” for the truthfulness concept. However, this strong association could be because the intervention and localization are “correlated”, since both use statistics of the same activations (determined by the design of the data, etc.). E.g. for heads with very low probing accuracy, the estimated intervention vectors could be very noisy, and thus the interventions could be less effective.
@@ 4 Finding “optimal” interventions
This formulation connects weight edits to representation edits. However, it doesn’t yet allow us to localize edits to specific heads — while $\bm{\theta}^{l}$ can be read as concatenation of headwise intervention vectors, the projected $W^{l}\bm{\theta}^{l}$ have no corresponding interpretations. Therefore, we can’t restrict the edits to specific heads by imposing structure on $\bm{b}^{l}$’s.
@@ 6 Intervening a single head is just as effective
It is now clear that edit-based evidence does not provide strong evidence for localization in the 16 head setup. However, a possible way of saving localization would be to argue that 16 heads is too many, giving too much leeway to induce any behavior we want with editing. For example, if we edited half the heads of the model, it would not be surprising if we could make the model do anything we wanted. Accordingly, we might hope that there is still a valid syllogism of the form “the localized edit is extremely constrained” and “edits at this location optimally control the target behavior” implies “the target behavior is localized to this location”.
To test this, we now focus on the single head case. The procedure is simple: we randomly sample $24$ single heads, one at a time, and search for optimal interventions. The distribution of the best Info*Truth scores is shown in fig. 4(a). We find $5$ single-heads that are as effective fig. 4(b), and none of them has high probing accuracy.
Notice that, still, none of these heads can be understood as localizing the truthfulness concept. The reason is that there are multiple distinct locations that work equally well! That is, even in the extreme case of a very localized edit that replicates the target behavior essentially optimally, we still cannot conclude that there is evidence supporting localization.
@@ 7 Are the Probing-Localized Heads Anything Special?
So far what we mean by localization, is that we can change model generation on target concept by an edit at this location. And our experiments show no evidence for this type of localization, and probing-localized heads play no special role.
@@
It’s important to note that the model’s probabilities for fixed responses, do not directly correspond to what the model actually generates. Even if the model assigns a higher probability to a truthful response than an untruthful one, it may still not generate the truthful response if the fixed dataset is off-policy (i.e. both probabilities are low). This highlights the well-known gap between what a model “knows” (which is the motivation behind probing) and what it ultimately generates [joshi2023personas, wang2020language, kadavath2022language, saunders2022self, burns2022discovering].